OpenStudy (anonymous):

y(y-3)=4, solve for y

OpenStudy (anonymous):

quad form. y^2-3y-4=0

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

y2-3y-4=0 y2-4y+y-4=0 y(y-4)+(y-4)=0 (y+1)(y-4)=0

OpenStudy (anonymous):

y^2-3y-4=0; then factor, we get (y-4)(y+1); values of y= 4, -1